Fx(x) L - - +--- - + -2 -1 1 2 CDF Ex(x)Let X have CDF Fx(2) shown in a) Find P[X > 0.5] and P[X > 1). b) Find E[X].
Answer #1

Answer:

Given that:

Fx(x) L - - +--- - + -2 -1 1 2 CDF Ex(x)F(x)=\left\{\begin{matrix} \frac{x}{4} & -2<x<1\\ \frac{1}{3} & x=1\\ \frac{x}{4}& 1<x<2\\ 1 & x\geq 2 \end{matrix}\right.

F:0) = 1: -2 <r <1 1<x< 2

F(x)=\left\{\begin{matrix} \frac{1}{4}; & -2<x<2\\ 0;& otherwise\\ \end{matrix}\right.

a) Find P[X>0.5] and P[X>1]

P(X>0.5)=\int_{0.5}^{2}f(x)dx

P(X>0.5)=\frac{1}{4}[2-0.5]

P(X>0.5)=\frac{1.5}{4}

P(X>1)=\int_{1}^{2}\frac{1}{4}dx

P(X>1)=\frac{1}{4}[2-1]

P(X>1)=\frac{1}{4}

P(X>1)=0.25

b) Find E[X]

E(X)=\int_{-2}^{2}x\, \frac{1}{4}dx

E(X)=\frac{1}{4}[\frac{x^2}{2}]_{-2}^{2}

E(X)=\frac{1}{8}[4-4]

E(X)=\frac{1}{8}[0]

E(X)=0